An evacuated glass vessel weghs 50.0 g when empty, 148.0 g when filled with a liquid of density 0.98 g mL-1 and 50.0 g when filledwith an ideal gas at 760 mmHg at 300K. Determine the molar mass of the gas

Grade:upto college level

2 Answers

Navjyot Kalra
9 years ago
Weight of liquid = 148 – 50 =98 g
Volume of liquid = 98/0.98 = 100 ml = volume of vessel it means, vessel of 100 ml contain ideal gas at 760 mm Hg at 300 K
Weight of gas = 50.5 - 50 = 0.5g
Using, PV = nRT = w/m RT
760/760 * 100/1000 = 0.5/m * 0.082 *300 [n = 0.5/m]
∴ Molecular weight of gas (m) = 123
Biswas
5 years ago
Since the vessel weigh 50.5g so the water and gas weighs 98 and 0.5 g respectively
Similarly..from the density and mass of the liquid we calculate the volume of water which is equivalent to volume of vessel and gas as well....now using ideal gas law PV=nRT we find (n) mole then mole is equal to ratioof given mass and molar mass we find molar mass.
Its thats simple😊
Then...using ideal gas law..
PV=nRT we find mole(n) and now mole is equivalent to